# A Comprehensive Review of Few-shot Action Recognition

### v1

@@ Abstract
Few-shot action recognition aims to address the high cost and impracticality of manually labeling complex and variable video data in action recognition. It requires accurately classifying human actions in videos using only a few labeled examples per class. Compared to few-shot learning in image scenarios, few-shot action recognition is more challenging due to the intrinsic complexity of video data. Recognizing actions involves modeling intricate temporal sequences and extracting rich semantic information, which goes beyond mere human and object identification in each frame. Furthermore, the issue of intra-class variance becomes particularly pronounced with limited video samples, complicating the learning of representative features for novel action categories. To overcome these challenges, numerous approaches have driven significant advancements in few-shot action recognition, which underscores the need for a comprehensive survey. Unlike early surveys that focus on few-shot image or text classification, we deeply consider the unique challenges of few-shot action recognition. In this survey, we review a wide variety of recent methods and summarize the general framework. Additionally, the survey presents the commonly used benchmarks and discusses relevant advanced topics and promising future directions. We hope this survey can serve as a valuable resource for researchers, offering essential guidance to newcomers and stimulating seasoned researchers with fresh insights.
Index Terms:
@@
I Introduction
@@
With the rapid advancements and innovations in communication and media creation technologies, video has increasingly emerged as a prominent form of information dissemination. Recognizing and understanding human actions from videos is a critical topic for numerous real-world applications, including traffic monitoring [1], sports competitions [2], and film production [3]. The emergence of deep learning has significantly advanced the frontiers of action recognition [4, 5, 6]. However, traditional deep learning models typically require a large amount of labeled video data to achieve optimal performance. This hampers the widespread application of deep learning techniques in many real-world scenarios, where obtaining and annotating extensive video datasets is both costly and time-prohibitive. Few-shot action recognition (FSAR) [7] has been proposed to address this challenge. FSAR aims to enable models to learn and generalize effectively with only a few labeled videos, thereby reducing data dependency and enhancing application efficiency.
@@
In contrast to few-shot learning in image or text scenarios [9, 10, 11], few-shot action recognition presents unique challenges because videos encompass both visual information within frames and temporal dynamics between frames. Recognizing human actions necessitates the careful modeling of temporal sequences and rich semantic information, beyond merely identifying human appearance, objects and the context in each frame. For instance, the action of “underwater diving” requires not only identifying water environments and divers but also understanding the trajectories to distinguish fine-grained diving actions such as “ascent”, “balance”, and “descent”, which are temporally distinct despite their similar appearances. This poses a significant challenge in extracting effective features for human action recognition. Consequently, learning intricate video representations that can be generalized to novel categories remains a difficult task. Moreover, the issue of intra-class variance becomes increasingly pronounced when only a few video samples are available. For example, videos of “underwater diving” can exhibit significant intra-class variance from both static and dynamic perspectives. Static variability includes different types of diving equipment and diverse environments. Dynamic variability encompasses variations in motion speed, attitude adjustment, and technical style. The few-shot video data with significant individual differences are always far from enough to describe the real distribution of the action. In this case, the model may become overly specialized to the limited videos, resulting in overfitting and failing to generalize effectively to novel instances of the same action. In summary, the complexity of videos and the inevitable sample-wise variance within the same category make it challenging to learn novel action concepts with only a few labeled samples.
@@
With the rapid advancement of FSAR, a substantial body of new literature is being produced to address the unique challenges. Figure 1 displays the performance trends of state-of-the-art methods over the years. Although recent relevant surveys [12, 13, 14, 15] have reviewed existing few-shot learning methods, they always overlook the unique characteristics of FSAR and cannot provide exhaustive summaries of recent research and findings in this field. This motivates us to write a survey that reviews various few-shot action recognition methods from a comprehensive and systematic perspective, synthesizing the latest research developments as a guide for researchers and practitioners. Unlike early surveys that either briefly mention few-shot action recognition as a future trend or lack comprehensive content coverage, we deeply consider the unique challenges and provide an exhaustive analysis of all recent methods.
@@
Reviewing the development of few-shot action recognition, early approaches are Generative-based which focus on generating additional data to mitigate sample scarcity [16, 17, 18, 19]. In contrast, the majority of recent studies have embraced meta-learning frameworks, emphasizing critical stages including Video Instance Representation, Category Prototype Learning, and Generalized Video Alignment. Figure 2 depicts a chronological overview of the development of different methods for few-shot action recognition. Building upon these foundations, we systematically categorize existing works and develop a novel taxonomy (as shown in Figure 3) to explore the contributions of existing methods, providing a thorough comparative analysis and discussion of recent advancements. Alongside this taxonomy, where existing FSAR approaches are categorized according to their contributions to addressing the challenges, we conduct a detailed analysis of the characteristics of existing approaches. Next, we highlight recent advancements in FSAR research within advanced fields, showcasing the methods developed to tackle application challenges. Finally, drawing from our comprehensive analysis, we propose promising research directions for future investigations. These insights are intended to inspire and guide future research endeavors, ultimately driving continued progress in the field.
Figure 2: A chronological overview of few-shot action recognition methods in different categories.
@@
Organization. In this paper, we first review the related surveys in Section II. Next, we examine the pertinent research areas, formulate the problem definition and the evaluation metrics, and introduce the benchmarks in Section III. We then review the few-shot action recognition methods in Section IV. Here, we organize previous methods into four types: generative-based methods, video instance representation learning, category prototype learning, and generalized video alignment. In Section V, we discuss the latest developments in few-shot action recognition topics. Section VI focuses on future research opportunities and discusses open issues. Finally, Section VII concludes our paper. Our survey aims to contribute to the advancement of few-shot action recognition while providing valuable insights to the scholarly community.
II Related Surveys
@@
Few-shot Action Recognition. Wang et al. [23] provide a brief overview of few-shot action recognition, yet their review lacks thorough comparative analysis, comprehensive content coverage, and recent advancement discussion.
Zero-shot Action Recognition. Estevam et al. [24] present a survey of zero-shot action recognition methods that describes several techniques used to perform visual and semantic extraction. They do not take into account the few-shot action recognition, as zero-shot learning is inherently different from the few-shot setting. In zero-shot learning, predictions are made based on semantic information rather than on a few samples.
@@ III Background
III-A Relevant Research Areas
@@
III-A1 Human action recognition
@@
Action recognition is a critical task in computer vision that involves identifying and categorizing human actions in video sequences. As illustrated in Figure 4, the task requires recognizing the categories of new videos by learning action concepts from a large number of labeled videos. Unlike static image classification, human action recognition must account for the temporal dynamics and sequential nature of actions, which significantly increases the complexity of the task. Common attempts leverage various advanced methods to accurately identify activities in video sequences, such as 3D Convolutional Neural Networks (CNNs) [27, 28, 29, 30, 31], RNN/LSTM networks [32, 33, 34], Attention mechanisms [35, 36], Graph-based methods [37, 38], Transformers [39, 40, 41], Self-supervised and unsupervised learning techniques [42, 43]. In recent years, the emergence of various sensors has enabled the representation of human actions through diverse modalities beyond RGB, including depth [44, 45], skeleton [46, 47], infrared sequences [48], point clouds [49], event streams [50], audio [51], acceleration [52, 53], radar [54], and WiFi [55]. Human action recognition has a wide range of applications and therefore has been attracting increasing attention in the field of computer vision, including surveillance [56], human-computer interaction [57], and sports analytics [2]. Despite significant progress in efficient and accurate action recognition, further exploration of advanced few-shot action recognition is still required. Different from conventional supervised action recognition tasks, few-shot action recognition necessitates models to generalize to novel actions using only a limited number of videos.
III-A2 Few-shot Learning
@@
Few-shot action recognition is a critical application of few-shot learning due to the high cost of obtaining sufficient labeled videos. Specifically, Few-shot learning (FSL) is a technique that focuses on training models that can generalize effectively from a very limited number of labeled training examples. In contrast to traditional data-driven machine learning paradigms, FSL aims to enable models to perform well on new tasks with only a few instances per class by leveraging prior knowledge and inductive biases learned from related tasks or datasets. It is particularly valuable in scenarios where data collection and annotation are expensive or impractical. We mainly talk about the few-shot classification tasks such as image classification [58, 59], sentiment classification [60], graph classification [61], and audio classification [62]. Few-shot classification often employs an $N$-way $K$-shot framework, where $N$ represents the number of classes and $K$ denotes the number of examples per class. The widely used meta-learning paradigm involves two phases: meta-training and meta-test. During meta-training, models are exposed to multiple episodes to learn a strategy for quick adaptation. During the meta-test, the model applies this strategy to new tasks with limited data. There are two types of meta-learning, including optimization-based methods [63, 64], which adjust the model parameters to facilitate fast learning on new tasks, and metric-based methods [58, 59, 65], which focus on learning a distance metric to compare new examples with stored representations of known classes. Additionally, other methods leverage data augmentation techniques to mitigate the issue of limited supervised information [66, 67]. These techniques involve generating synthetic data that augment the existing small dataset, thereby providing more examples for the model to learn from. Compared to conventional few-shot classification tasks, few-shot action recognition is more challenging due to the additional temporal domain information in video data.
III-B Problem Definition
@@
1 for $\mathcal{T}^{(e)}$=$\mathcal{T}_{1},\dots,\mathcal{T}^{(E)}$ do
2       Randomly select $N$ novel classes.
3       Randomly select $K$ samples from each class to construct the support set $\mathcal{S}^{(e)}=\{(x_{i},y_{i})\}_{{}_{i=1}}^{{}^{NK}}$.
4       Randomly select $M$ samples from the remaining samples of $N$ classes as the query set $\mathcal{Q}^{(e)}=\{x_{i}\}_{{}_{i=1}}^{{}^{M}}$, with the labels $Y=\{y_{i}\}_{{}_{i=1}}^{{}^{M}}$.
5       Record predicted labels $\hat{Y}^{(e)}=\phi(\mathcal{Q},\mathcal{S})$.
6       Compute accuracy $a^{(e)}=\frac{1}{M}\sum_{j=1}^{M}\mathbb{1}[\hat{y}_{j}^{(e)}=y_{j}^{(e)}]$.
7
8 end for
@@
Few-shot action recognition refers to the task of effectively identifying human actions in unlabeled videos using a limited number of labeled samples from novel action categories (Figure 4). The task contains a base dataset $\mathbb{D}_{train}$, which includes ample labeled examples from base action classes $C_{train}$, and a novel dataset $\mathbb{D}_{test}$ from novel classes $C_{test}$ to assess the model’s performance on previously unseen classes, ensuring $C_{test}\cap C_{train}=\varnothing$. In the evaluation phase, the $N$-way $K$-shot setting involves novel $N$ classes and $K$ examples from each class to form a support set, while an additional query set is used to evaluate the model’s ability to generalize from the limited examples. As shown in Algorithm 1, the evaluation procedure includes many episodes. Each episode is comprised of a query set $\mathcal{Q}\subset\mathbb{D}_{test}$ and a support set $\mathcal{S}\subset\mathbb{D}_{test}$. The query set $\mathcal{Q}=\{(x_{i},y_{i})\}_{{}_{i=1}}^{{}^{M}}$ contains $M$ query samples, where $x_{i}$ denotes the video of the $i^{th}$ query sample $x_{i}$, and $y_{i}\in\{1,2,...,N\}$ denotes the class label. The support set $\mathcal{S}=\{(x_{i},y_{i})\}_{{}_{i=1}}^{{}^{NK}}$ contains ${K}$ samples for each of the $N$ classes. Note that the class label of each query sample is invisible during the meta-test. The performance of the learning algorithm is measured by the classification accuracy averaged over all episodes.
III-C Benchmarks
TABLE I: Benchmarks for few-shot action recognition.

### v2

@@ Abstract
Few-shot action recognition aims to address the high cost and impracticality of manually labeling complex and variable video data in action recognition. It requires accurately classifying human actions in videos using only a few labeled examples per class. Compared to few-shot learning in image scenarios, few-shot action recognition is more challenging due to the intrinsic complexity of video data. Numerous approaches have driven significant advancements in few-shot action recognition, which underscores the need for a comprehensive survey. Unlike early surveys that focus on few-shot image or text classification, we deeply consider the unique challenges of few-shot action recognition. In this survey, we provide a comprehensive review of recent methods and introduce a novel and systematic taxonomy of existing approaches, accompanied by a detailed analysis. We categorize the methods into generative-based and meta-learning frameworks, and further elaborate on the methods within the meta-learning framework, covering aspects: video instance representation, category prototype learning, and generalized video alignment. Additionally, the survey presents the commonly used benchmarks and discusses relevant advanced topics and promising future directions. We hope this survey can serve as a valuable resource for researchers, offering essential guidance to newcomers and stimulating seasoned researchers with fresh insights.
keywords:
@@
1 Introduction
Figure 1: A taxonomically organized chronological timeline of few-shot action recognition methodologies, highlighting key developments and emerging trends. The plot illustrates the field’s growth by showcasing the cumulative temporal distribution of publications from 2018 to 2025. The red underline signifies the method employs a language–image pretraining backbone (e.g., CLIP), while others adopt a unimodal pretraining backbone (e.g., ResNet).
@@ 1 Introduction
With the rapid advancements and innovations in communication and media creation technologies, video has increasingly emerged as a prominent form of information dissemination. Recognizing and understanding human actions from videos is a critical topic for numerous real-world applications. The emergence of deep learning has significantly advanced the frontiers of action recognition [152, 153]. However, traditional deep learning models typically require a large amount of labeled video data to achieve optimal performance. This hampers the widespread application of deep learning techniques in many real-world scenarios, where obtaining and annotating extensive video datasets is both costly and time-prohibitive. Few-shot action recognition (FSAR) [196] has been proposed to address this challenge. FSAR aims to enable models to learn and generalize effectively with only a few labeled videos, thereby reducing data dependency and enhancing application efficiency.
@@
In contrast to few-shot learning in image or text scenarios [106, 9, 115], few-shot action recognition presents unique challenges because videos encompass both visual information within frames and temporal dynamics between frames. Recognizing human actions necessitates the careful modeling of temporal sequences and rich semantic information, beyond merely identifying human appearance, objects, and the context in each frame. For instance, the action of “underwater diving” requires not only identifying water environments and divers but also understanding the trajectories to distinguish fine-grained diving actions such as “ascent”, “balance”, and “descent”, which are temporally distinct despite their similar appearances. This poses a significant challenge in extracting effective features for human action recognition. Consequently, learning intricate video representations that can be generalized to novel categories remains a difficult task. Moreover, the issue of intra-class variance becomes increasingly pronounced when only a few video samples are available. For example, videos of “underwater diving” can exhibit significant intra-class variance from both static and dynamic perspectives. Static variability includes different types of diving equipment and diverse environments. Dynamic variability encompasses variations in motion speed, attitude adjustment, and technical style. The few-shot video data with significant individual differences are always far from enough to describe the real distribution of the action. In this case, the model may become overly specialized to the limited videos, resulting in overfitting and failing to generalize effectively to novel instances of the same action. In summary, the complexity of videos and the inevitable sample-wise variance within the same category make it challenging to learn novel action concepts with only a few labeled samples.
@@
With the rapid advancement of FSAR, a substantial body of new literature is being produced to address the unique challenges. Although recent relevant surveys [163, 72, 36, 123] have reviewed existing few-shot learning methods, they always overlook the unique characteristics of FSAR and cannot provide exhaustive summaries of recent research and findings in this field. This motivates us to write a survey that reviews various few-shot action recognition methods from a comprehensive and systematic perspective, synthesizing the latest research developments as a guide for researchers and practitioners. Unlike early surveys that either briefly mention few-shot action recognition as a future trend or lack comprehensive content coverage, we deeply consider the unique challenges and provide an exhaustive analysis of all recent methods.
Reviewing the development of few-shot action recognition, early approaches are Generative-based which focus on generating additional data to mitigate sample scarcity [96, 66, 30, 174]. In contrast, the majority of recent studies have embraced meta-learning frameworks, emphasizing critical stages including Video Instance Representation, Category Prototype Learning, and Generalized Video Alignment. Figure 1 depicts a chronological overview of the development of different methods for few-shot action recognition. As illustrated, there is a gradual rise in the quantity of FSAR publications, and generative-based methods are concentrated in the earlier period, whereas the subsequent techniques are primarily rooted in the meta-learning framework. It is also notable that methods based on multimodal pre-trained backbones emerged in 2023.
@@
We systematically categorize existing works and develop a novel taxonomy (as shown in Figure 2) to explore the contributions of existing methods, providing a thorough comparative analysis and discussion of recent advancements. Alongside this taxonomy, where existing FSAR approaches are categorized according to their contributions to addressing the challenges, we conduct a detailed analysis of the characteristics of existing approaches. Next, we highlight recent advancements in FSAR research within advanced fields, showcasing the methods developed to tackle application challenges. Finally, drawing from our comprehensive analysis, we propose promising research directions for future investigations. These insights are intended to inspire and guide future research endeavors, ultimately driving continued progress in the field.
@@
Organization. In this paper, we first review the related surveys in Section 2. Next, we examine the pertinent research areas, formulate the problem definition and the evaluation metrics, and introduce the benchmarks in Section 3. We then review the few-shot action recognition methods in Section 4. Here, we organize previous methods into four types: generative-based methods, video instance representation learning, category prototype learning, and generalized video alignment. In Section 5, we discuss the latest developments in few-shot action recognition topics. Section 6 focuses on future research opportunities and discusses open issues. Finally, Section 7 concludes our paper. Our survey aims to contribute to the advancement of few-shot action recognition while providing valuable insights to the scholarly community.
2 Related Surveys
@@
Few-shot Action Recognition. Wang et al. [146] and Ruan et al. [110] provide brief overviews of few-shot action recognition. Specifically, Wang et al. [146] summarized FSAR methods according to the task process, and Ruan et al. [110] simply categorize FSAR methods into: memory networks, metric learning, and data augmentation. While these taxonomies bear a resemblance to general FSL surveys [72, 36], they do not specifically highlight the contributions of current methodologies in addressing the unique challenges within FSAR. And these surveys fail to encompass all significant works since the establishment of this field and are limited in scope, lacking comprehensive comparative analysis, detailed taxonomy, and discussion of recent advancements. On the contrary, we present the most systematic review in the field of FSAR, extending to advanced topics and broadening its applicability to more intricate contexts. It includes valuable details and fills the gaps left by existing surveys. Zero-shot Action Recognition. Estevam et al. [25] present a survey of zero-shot action recognition methods that describes several techniques used to perform visual and semantic extraction. They do not take into account the few-shot action recognition, as zero-shot learning is inherently different from the few-shot setting. In zero-shot learning, predictions are made based on semantic information rather than on a few samples.
@@ 3 Background
3.1 Relevant Research Areas
@@
3.1.1 Human action recognition
@@
Action recognition is a critical task in computer vision that involves identifying and categorizing human actions in video sequences. As illustrated in Figure 3, the task requires recognizing the categories of new videos by learning action concepts from a large number of labeled videos. Unlike static image classification, human action recognition must account for the temporal dynamics and sequential nature of actions, which significantly increases the complexity of the task. Common attempts leverage various advanced methods to accurately identify activities in video sequences, such as 3D Convolutional Neural Networks (CNNs) [56, 132, 137], RNN/LSTM networks [43], Attention mechanisms [69], Graph-based methods [119], Transformers [85], Self-supervised and unsupervised learning techniques [145]. In recent years, the emergence of various sensors has enabled the representation of human actions through diverse modalities beyond RGB, including depth [109], skeleton [149, 142], infrared sequences [2], point clouds [95], event streams [55], audio [75], acceleration [5], radar [76], and WiFi [199]. Human action recognition has a wide range of applications and therefore has been attracting increasing attention in the field of computer vision, including surveillance [49], human-computer interaction [4], and sports analytics [124]. Despite significant progress in efficient and accurate action recognition, further exploration of advanced few-shot action recognition is still required. Different from conventional supervised action recognition tasks, few-shot action recognition necessitates models to generalize to novel actions using only a limited number of videos.
3.1.2 Few-shot Learning
@@
Few-shot action recognition is a critical application of few-shot learning due to the high cost of obtaining sufficient labeled videos. Specifically, Few-shot learning (FSL) is a technique that focuses on training models that can generalize effectively from a very limited number of labeled training examples. In contrast to traditional data-driven machine learning paradigms, FSL aims to enable models to perform well on new tasks with only a few instances per class by leveraging prior knowledge and inductive biases learned from related tasks or datasets. It is particularly valuable in scenarios where data collection and annotation are expensive or impractical. We mainly talk about the few-shot classification tasks such as image classification [144, 122], sentiment classification [186], graph classification [88], and audio classification [164].
Few-shot classification often employs an $N$-way $K$-shot framework, where $N$ represents the number of classes and $K$ denotes the number of examples per class. In practical application, there are more universal settings: Generalized FSL [174] encompasses the capability of models to recognize both few-shot categories and seen categories, reflecting real-world complexity. Open-set FSL [8, 77] aims to recognize few-shot classes and reject samples that do not belong to either seen or few-shot classes, with the goal of extending open-set recognition beyond the large-scale setting. Open-world FSL [116, 3] is a recent research field dedicated to accurately identifying target samples in scenarios where data is limited and labels are unreliable. This research holds significant practical implications and is highly relevant to real-world applications. The main challenge of these tasks is the same as the conventional few-shot setting: learning a new category with a few samples and identifying its category when given a new sample. Therefore, researching prevalent FSL algorithms is valuable and necessary.
Prominent FSL methodologies encompass meta-learning methods, data augmentation techniques, amortization-based methods, and in-context learning methods. Meta-learning is a widely used paradigm in FSL, which involves two phases: meta-training and meta-test. During meta-training, models are exposed to multiple episodes to learn a strategy for quick adaptation. During the meta-test, the model applies this strategy to new tasks with limited data. There are two types of meta-learning, including optimization-based methods [29, 84], which adjust the model parameters to facilitate fast learning on new tasks, and metric-based methods [144, 122, 120], which focus on learning a distance metric to compare new examples with stored representations of known classes. Several other methods leverage data augmentation techniques to mitigate the issue of limited supervised information [192, 184]. These techniques involve generating synthetic data that augment the existing small dataset, thereby providing more examples for the model to learn from. Additionally, amortization-based methods (e.g. CNPs [35]) combine the advantages of deep neural networks and Bayesian methods, enabling accurate predictions by observing only a handful of training data points. Recently, with the advancement of Large Language Models (LLMs), in-context learning (ICL) has been proposed to address FSL [51]. In this framework, an LLM receives a test instance and a few examples as input and directly performs recognition or reasoning without requiring any parameter updates. Compared to conventional few-shot classification tasks, few-shot action recognition is more challenging due to the additional temporal domain information in video data.
@@
for $\mathcal{T}^{(e)}$=$\mathcal{T}^{(1)},\dots,\mathcal{T}^{(E)}$ do
       Randomly select $N$ novel classes.
       Randomly select $K$ samples from each class to construct the support set $\mathcal{S}^{(e)}=\{(x_{i},y_{i})\}_{{}_{i=1}}^{{}^{NK}}$.
       Randomly select $M$ samples from the remaining samples of $N$ classes as the query set $\mathcal{Q}^{(e)}=\{x_{i}\}_{{}_{i=1}}^{{}^{M}}$, with the labels $Y=\{y_{i}\}_{{}_{i=1}}^{{}^{M}}$.
       Record predicted labels $\quad\quad\hat{Y}^{(e)}=\phi(\mathcal{Q},\mathcal{S})$.
       Compute accuracy
       $\quad\quad a^{(e)}=\frac{1}{M}\sum_{j=1}^{M}$ $\mathds{1}$ $[\hat{y}_{j}^{(e)}=y_{j}^{(e)}]$.
end for
@@
3.2 Problem Definition
Few-shot action recognition refers to the task of effectively identifying human actions in unlabeled videos using a limited number of labeled samples from novel action categories (Figure 3). The task contains a base dataset $\mathbb{D}_{train}$, which includes ample labeled examples from base action classes $C_{train}$, and a novel dataset $\mathbb{D}_{test}$ from novel classes $C_{test}$ to assess the model’s performance on previously unseen classes, ensuring $C_{test}\cap C_{train}=\varnothing$. In the evaluation phase, the $N$-way $K$-shot setting involves novel $N$ classes and $K$ examples from each class to form a support set, while an additional query set is used to evaluate the model’s ability to generalize from the limited examples. As shown in Algorithm 1, the evaluation procedure includes many episodes. Each episode ($\mathcal{T}$) is comprised of a query set $\mathcal{Q}\subset\mathbb{D}_{test}$ and a support set $\mathcal{S}\subset\mathbb{D}_{test}$. The query set $\mathcal{Q}=\{(x_{i},y_{i})\}_{{}_{i=1}}^{{}^{M}}$ contains $M$ query samples, where $x_{i}$ denotes the video of the $i^{th}$ query sample $x_{i}$, and $y_{i}\in\{1,2,...,N\}$ denotes the class label. The support set $\mathcal{S}=\{(x_{i},y_{i})\}_{{}_{i=1}}^{{}^{NK}}$ contains ${K}$ samples for each of the $N$ classes. Note that the class label of each query sample is invisible during the meta-test. The performance of the learning algorithm is measured by the classification accuracy averaged over all episodes.
3.3 Benchmarks
Table 1: Benchmarks for few-shot action recognition.